OpenStudy (anonymous):

!PLEASE HELP ME WITH THIS PROBLEM SOMEBODY:(! A triangle paddock has a road (AB)forming its longest side.AB is 13 km long.The AC and CB are at right angels. if BC is 400 m longer than AC.Find the area of the paddock in hectares?

OpenStudy (campbell_st):

so let AC be length x then CB is 0.4 km longer so CB = 1.4x  so using pythagoras's theorem you will have \[13^2 = x^2 + (1.4x)^2\] now just solve for x.

OpenStudy (anonymous):

SO MY ANSWERS WILL BE X=-1.74371 AND +1074371

OpenStudy (kamille):

x can't be with "-" in geometry probelms. Why?

OpenStudy (campbell_st):

well only consider the positive answer... as measurements can't be negative so so x is 7.556 km... or 7556 metres... so the base of the triangle is 7556 metres and the height is 7956 metres so to find the area \[A = \frac{1}{2} \times 7556 \times 7956 \] to get the number of hectares.... use the conversion 1ha = 100 x 100 or 1ha = 10000 m^2

OpenStudy (anonymous):

SO THEN IT EQUALS A=72386432

OpenStudy (campbell_st):

well I got the area as A = 30057768 m^2 divide by 10000 to get the hectacres

OpenStudy (anonymous):

A=3006M^2

OpenStudy (campbell_st):

well check the units.... ha... not m^2... but it seems to make sense...

OpenStudy (anonymous):

so it wouldn't be on ^2

OpenStudy (campbell_st):

no the question asked for hectacres... where are just ha.. they are a measurement of area..